# 3D Kinematics Estimation from Video with a Biomechanical Model and Synthetic Training Data

### 26 Feb 2024 v2

@@ 1 Introduction
Human motion capture is the process of capturing the relative position of human body segments through time. In biomechanics, accurate estimation of joint loading of the human body is crucial for applications in healthcare and sports, as it helps in understanding the biomechanical stresses experienced by the joints during movement. Biomechanical models can provide these estimates but to be clinically successful, these models must accurately describe the musculoskeletal system and have accurate estimates of the joint kinematics. Optoelectronic motion capture systems are the established standard in human motion capture [30]. These systems necessitate the attachment of reflective markers to the bony landmarks of the subject. As a result, this method is time-consuming, requires a substantial financial investment and the markers also restrict the natural movement of the individuals [7, 18]. To address these issues, researchers have focused on developing markerless motion capture methods that utilize visual inputs and machine learning techniques to estimate human motion without the need for physical markers and specialized expertise [21, 29, 1, 27].
Most of the existing markerless motion capture methods utilize 2D-to-3D lifting 3D human pose estimation methods, such as OpenPose [2], as the backbone to infer 3D poses from detected 2D joints. These conventional methods usually take a two-step approach that first detects human body joint with an off-the-shelf tool [2] and then performs inverse kinematic to retrieve the skeleton of the body [29, 21]. Therefore, the estimation accuracy of these methods relies on the body joint detection in the first step. The end-to-end attempt has been proposed for the task with deep learning techniques, which outperform the previous two-step approach and show a promising direction [1]. However, the end-to-end approach has not been fully explored yet for the 3D kinematic estimation task. On another hand, 3D kinematics estimation of the human body requires accurate anatomic definition of joints. The gold standard way to measure human joint kinematics is through the X-ray scan of the person [33, 11], which is not realistic to conduct in daily life settings. As an alternative, marker-based motion capture is incorporated to create datasets with accurate 3D kinematics annotations [26, 29, 8]. Still, the expensive and time-consuming nature of marker-based motion capture impedes the construction of large-scale datasets necessary for deep-learning-based markerless motion capture networks [1, 20]. Last, many 3D human pose estimation methods overlook the incorporation of biomechanical constraints, leading to predictions that deviate from anatomically realistic movements. This discrepancy is especially problematic for 3D kinematics estimation, which demands a higher level of biomechanical fidelity.
@@
In this work, we propose a novel markerless motion capture framework comprising a biomechanics-aware network, as illustrated in Figure 1. The model leverages two-view RGB videos as inputs and performs feature extraction [5], followed by a feature aggregation step to generate frame features. There is no explicit human body joint detection, instead, we sample the feature point on the input frame. Given the dynamic nature of human motion, the frame features are further refined by integrating temporal information through a transformer-based U-Net architecture. Finally, the model outputs a sequence of joint angles and a set of body segment scales.
@@ 2.1 Markerless Motion Capture with OpenSim
Markerless motion capture draws people’s attention with its cost efficient nature and its ability to yield comparable results to marker-based motion capturing [27, 4]. Advancements in 3D human motion estimation have enabled the integration of 3D human motion estimation techniques with biomechanical models, allowing for a comprehensive analysis of human biomechanics. Most of the existing markerless motion capture methods use multi-step processing for kinematics estimation [21, 29]. The process starts with deriving 3D joint positions by triangulating the detected 2D landmarks [2] from multiple views. Next, the 3D joint positions are treated as marker positions in the Inverse Kinematics (IK) tool and scaling tool in OpenSim software to derive joint kinematics and body segment scales. The mapping from the detected 3D joints to the real 3D joints can be encoded in the marker offsets defined in the OpenSim model [21]. Alternatively, it is possible to train a model to learn this mapping function [29]. To further improve the performance, the landmark confidence scores are considered to remove low-confidence landmarks [21, 29].
@@
With the direct human pose estimation strategy, the proposed method uses a frame feature encoder to map the input frame into the frame features with sampling on the input video frames. Different from previous works [5], we largely improve the architecture to work with two input views to perform 3D kinematic estimation. In addition, we propose a spatio-temporal refinement module with a transformer-based U-Net architecture adapted from [34], which was originally proposed for image denoise. The combination of transformer and U-Net facilitates temporal feature learning across different temporal resolutions while concurrently refining local spatial features.
@@ 3.1 Overview
To effective 3D kinematics estimation, the proposed end-to-end biomechanics-aware network first encodes each frame, captured from two different views, into a frame feature using a stacked hourglass network [19] and multilayer perceptron (MLP). Once the frame features of the entire sequence are obtained, a spatio-temporal feature refinement is performed to jointly consider spatial and temporal information to optimize the features across frames. To achieve accurate kinematics annotations, we create a synthetic dataset ODAH with the combination of skeleton from OpenSim [6, 25] model, meshes from SMPL-X [22], and acquire the body shape and motion for the mesh from the subject-specific OpenSim skeletal model and joint angles.
@@ 3.4 Synthetic Data Generation
Skeleton and mesh registration. An OpenSim model is composed of rigid bodies (bones) that are connected by joints. Joints connect two reference frames: one on the parent and one on the child body of the joint, which coincide in space at the joint center. The SMPL-X and OpenSim skeletal model do not share the same joint definitions. The OpenSim skeletal model is based on human anatomical structures, whereas the armature of the SMPL-X model is based on matching the mesh to surface scans [11] instead of skeletal anatomy. To remove this discrepancy, we trained a joint regressor for the SMPL-X model that located the joint keypoints corresponding to the OpenSim model joint centers based on the position of the SMPL-X vertices. A variety of poses are included during the training of the joint regressor to avoid overfitting.
@@ 4.4 Baselines
For Pose2Sim and OpenCap, frontal and sagittal views are taken as inputs and the Body25 model in OpenPose [2] is used as the 2D landmark detection backbone. We use Pose2Sim default Butterworth low-pass filter with a cut-off frequency of six Hz as the smoothing filter. For OpenCap, the cut-off frequency is set to half of the framerate. To ensure a fair comparison, we disable the landmark synchronization and the video trimming in OpenCap. For D3KE, only the frontal view is utilized as the input, and we choose the transformer-based temporal model with the sequence length set to 243 frames. Additionally, we only compare joint angles present in our generic OpenSim skeletal model and exclude arm flexion since D3KE’s generic OpenSim skeletal model does not have a joint angle defined for arm flexion for the sake of fair comparison.

### 05 Mar 2024 v4

@@ 1 Introduction
Kinematics estimation is the process of capturing the relative position of human body segments through time. In biomechanics, accurate estimation of joint loading of the human body is crucial for applications in healthcare and sports, as it helps in understanding the biomechanical stresses experienced by the joints during movement. Optoelectronic motion capture systems are the established standard in Kinematics estimation [32]. These systems necessitate the attachment of reflective markers to the bony landmarks of the subject. As a result, this method is time-consuming, requires a substantial financial investment and the markers also restrict the natural movement of the individuals [8, 20]. Consequently, many researchers have focused on developing markerless kinematics estimation [23, 31, 1, 29], which usually utilize 3D human pose estimation methods, such as OpenPose [2], as the backbone to infer 3D poses from detected 2D joints. The recent development of deep learning methods has propelled substantial advancements in the field of 3D human pose estimation from detected 2D joints [3, 16, 38, 7]. It provides great potential through the integration of deep learning 3D human pose estimation techniques and biomechanics models.
Unfortunately, there are significant gaps between the two research fields. The primary issue is the 2D joint annotations for conventional human pose estimation [2] are often anatomically wrong, resulting in imprecise 3D kinematics estimation. Many 3D human pose estimation methods overlook the incorporation of biomechanical constraints, leading to predictions that deviate from anatomically realistic movements. This discrepancy is especially problematic for 3D kinematics estimation, which demands a higher level of biomechanical fidelity for clinical diagnosis and sports science domains.
Another challenge is the scarcity of large-scale datasets with accurate 3D biomechanics annotations. The gold standard way to measure human joint kinematics is through the X-ray scan of the person [35, 13], which is not realistic to conduct in daily life settings. Although some datasets incorporate maker-based motion capture systems [28, 31, 9], they suffer from sensor noise and lack synchronization with the recording videos. Additionally, the expensive and time-consuming nature of marker-based motion capturing further impedes the construction of large-scale datasets necessary for deep-learning-based markerless motion capture models [1].
@@
In this work, we propose a novel markerless motion capture framework comprising a biomechanics-aware network, as illustrated in Figure 1. The model leverages two-view RGB videos as inputs and performs feature extraction [5], followed by a feature aggregation step to generate frame features. There is no explicit human body joint detection, instead, we sample the feature point on the input frame. Given the dynamic nature of human motion, the frame features are further refined by integrating temporal information through a transformer-based U-Net architecture [36]. Finally, the model outputs a sequence of joint angles and a set of body segment scales.
@@ 2.1 Markerless Motion Capture with OpenSim
Markerless motion capture draws people’s attention with its cost efficient nature and its ability to yield comparable results to marker-based motion capturing [29, 4]. Advancements in 3D human motion estimation have enabled the integration of 3D human motion estimation techniques with biomechanical models, allowing for a comprehensive analysis of human biomechanics [12]. Most of the existing markerless motion capture methods use multi-step processing for kinematics estimation [23, 31]. The process starts with deriving 3D joint positions by triangulating the detected 2D landmarks [2] from multiple views. Next, the 3D joint positions are treated as marker positions in the Inverse Kinematics (IK) tool and scaling tool in OpenSim software to derive joint kinematics and body segment scales. The mapping from the detected 3D joints to the real 3D joints can be encoded in the marker offsets defined in the OpenSim model [23]. Alternatively, it is possible to train a model to learn this mapping function [31]. To further improve the performance, the landmark confidence scores are considered to remove low-confidence landmarks [23, 31].
@@
With the direct human pose estimation strategy, the proposed method uses a frame feature encoder to map the input frame into the frame features with sampling on the input video frames. Different from previous works [5], we largely improve the architecture to work with two input views to perform 3D kinematic estimation. In addition, we propose a spatio-temporal refinement module with a transformer-based U-Net architecture adapted from [36], which was originally proposed for image denoise.
@@ 3.1 Overview
The proposed biomechanics-aware network first encodes each frame into a frame feature using a stacked hourglass network [21]. Similar to previous works [31, 23], we utilize two different views for each frame to provide an absolute 3D kinematic scale, which is essential for many clinical and sports science applications. Then a spatio-temporal feature refinement is performed to jointly consider spatial and temporal information to optimize the features across frames. The final output is the joint parameters of the OpenSim model [6]. To achieve accurate kinematics annotations, we create a synthetic dataset ODAH with the combination of skeleton from OpenSim [6, 27] model, meshes from SMPL-X [24], and acquire the body shape and motion for the mesh from the subject-specific OpenSim skeletal model and joint angles.
@@ 3.4 Synthetic Data Generation
Skeleton and mesh registration. An OpenSim model is composed of rigid bodies (bones) that are connected by joints. Joints connect two reference frames: one on the parent and one on the child body of the joint, which coincide in space at the joint center. The SMPL-X and OpenSim skeletal model do not share the same joint definitions. To remove this discrepancy, we trained a joint regressor for the SMPL-X model that located the joint keypoints corresponding to the OpenSim model joint centers based on the position of the SMPL-X vertices. A variety of poses are included during the training of the joint regressor to avoid overfitting.
@@ 4.4 Baselines
For Pose2Sim and OpenCap, frontal and sagittal views are taken as inputs and the Body25 model in OpenPose [2] is used as the 2D landmark detection backbone. We use Pose2Sim default Butterworth low-pass filter with a cut-off frequency of six Hz as the smoothing filter. For OpenCap, the cut-off frequency is set to half of the framerate. To ensure a fair comparison, we disable the landmark synchronization and the video trimming in OpenCap. For D3KE, it only requires a single-view image as input that we take frontal view. We acknowledge that it is not a totally fair comparison due to different inputs. However, it is natural of the D3KE method with only single-view as the input. We choose the transformer-based temporal model with the sequence length set to 243 frames for the D3KE method. Additionally, we only compare joint angles present in our generic OpenSim skeletal model and exclude arm flexion since D3KE’s generic OpenSim skeletal model does not have a joint angle defined for arm flexion for the sake of fair comparison.